For the titration of 20.00 mL of 0.1206 M NH3 with 0.0947 M HCl, calculate the pH before the addition of any titrant, at 10.00 mL prior to equivalence point, at the equivalence point, and at 10.00 mL after the equivalence point. (Kb (NH3) = 1.76 x 10-5 )
1
Expert's answer
2020-05-28T10:33:15-0400

The reaction between ammonia and hydrochloric acid is the following

NH3*H2O+ HCl = NH4Cl + H2O


Initially, without addition of hydrochloric acid, the solution contains only ammonia.

The reaction of ammonia dissociation is the following

NH3*H2O <-> NH4+ + OH- with Kb = [NH4+]*[OH-]/[NH3*H2O] = 1.76*10-5

[NH4+] = [OH-] -> Kb = [OH-]2/[NH3*H2O];

[OH-] = Kw/[H+] -> Kb = (Kw/[H+])2/[NH3*H2O] -> [H+] = Kw/(Kb*[NH3*H2O])0.5 = 6.87*10-12 mol/L

pH = 11.16


In the equivalent point c(NH3*H2O)*V(NH3*H2O) = c(HCl)*V(HCl)

V(HCl) = 0.1206 mol/L*20 ml/0.0947 mol/L = 25.47 mL


At 10 ml prior to equivalence point (at 15.47 ml of added HCl solution) pH value of the solution is calculated as

pH = 14 - pKb + lg([NH3*H2O]/[NH4Cl]) = 14 - 4.75 - lg(([NH3*H2O]initial-[NH3*H2O]titrated)/[HCl]added) = 9.06


In the equivalence point

Ka = Kw/Kb

Let x mol/L of NH4+ be formed during neutralization of ammonia by HCl, x mol/L of NH3 will react and (0.1206 mol/L*20 ml/(20 ml + 25.47 ml)) = 0.053 mol/L is the initial amount of the NH3 in the solution. After the reaction with HCl it will remain (0.053 -x) mol/L of NH3

Then the balance between NH3 and NH4+ will be written using Ka (Ka = Kw/Kb)

In accordance with Ka = [NH3]*[H+]/[NH4+]

x2/(0.053-x) = 10-14/(1.76*10-5)

x = 5.49*10-6 mol/L

pH = -lgx = 5.26 in the equivalence point


After adding extra 10 ml of HCl (25.47+10 = 35.47 ml) after equivalence point

pH value is determined by the amount of HCl added

c(H+) = (10 ml*0.0947 mol/L)/35.47 ml = 0.0267 mol/L

pH = -lg(c(H+)) = 1.57